# A Unified Framework for Model Editing

### v2

@@ 6 Conclusion
In this paper we unite two popular model editing techniques, ROME and MEMIT, under the preservation-memorization objective, with ROME performing equality-constrained edits and MEMIT operating under a least-square constraint. We disentangle the edit-distribution algorithm proposed in MEMIT from the optimization objective, presenting them as separate entities, emphasizing that a fair comparison of future model editing techniques with MEMIT should be based on the objective of MEMIT rather than conflating it with the edit-distribution algorithm.
Finally, we present EMMET - Equality-constrained Mass Model Editing in a Transformer, a new batched-editing algorithm based on the preservation-memorization objective where batched-memorization happens under an equality constraint. Our experiments show that EMMET performs similarly to MEMIT across multiple dimensions and metrics. EMMET completes batched editing using both types of objectives and truly unifies model editing under the preservation memorization framework. We hope that this unifying framework improves the intuitive understanding of these algorithms and fuels future research based on both intuition and mathematics.
@@ 7 Limitations
While our technique may streamline error correction processes, it does not address deeper structural limitations within models, such as edited models inadvertently amplifying existing errors or introducing new inaccuracies. Furthermore, the effectiveness of our method varies depending on the complexity of the model architecture and the nature of the edited knowledge as evidenced by our experiments. Despite having a theoretically ‘stronger’ memorization objective, EMMET is not able to outperform MEMIT, which also indicates that we might have reached a saturation point for model editing using naive implementations of the preservation-memorization objective, underscoring the fact that significant progress is yet to be made in understanding edit distribution and its implications. Thus, while our work contributes to a deeper understanding of model behavior, it is essential to recognize and account for these limitations in the interpretation and application of our findings.
@@ A.1 Related Work
A.2 Implementation Details for ROME, MEMIT and EMMET
@@
A.3 Key-Value creation in ROME/MEMIT
@@
A.4 EMMET and MEMIT Downstream Performance Comparison

### v4

@@ 6 Conclusion
In this paper we unite two popular model editing techniques, ROME and MEMIT, under the preservation-memorization objective, with ROME performing equality-constrained edits and MEMIT operating under a least-square constraint. We disentangle the edit-distribution algorithm proposed in MEMIT from the optimization objective, presenting them as separate entities. We also present EMMET, a new batched-editing algorithm based on the preservation-memorization objective, where memorization happens under an equality constraint. Our experiments show that EMMET has similar performance to MEMIT across multiple dimensions and metrics.
Enabling batched editing with equality-constraint in the form of EMMET allows us to truly unify ROME and MEMIT and shows that both these algorithms are essentially equivalent in terms of their (i) optimization objective, (ii) their abilities (singular and batched editing, a symmetry enabled by EMMET), (iii) their model editing performance and (iv) their limitations (similar model degradation). EMMET is a cornerstone in completing this larger picture. These results suggest that EMMET (or ROME) and MEMIT not only have very similar theoretical roots but also perform similarly in practice. The unified framework presented in our work along with the disentanglement of edit distribution algorithm has also enabled a fair comparison between the two algorithms, which was not possible before our work. We hope that this framework facilitates ease of comparison, consistency of implementation, and a much deeper understanding of these model editing methods.
@@ 7 Limitations
While our technique may streamline error correction processes, it does not address deeper structural limitations within models, such as edited models inadvertently amplifying existing errors or introducing new inaccuracies. Furthermore, the effectiveness of our method varies depending on the complexity of the model architecture and the nature of the edited knowledge as evidenced by our experiments. Despite having a theoretically ‘stronger’ memorization objective, EMMET is not able to outperform MEMIT, which also indicates that we might have reached a saturation point for model editing using naive implementations of the preservation-memorization objective, underscoring the fact that significant progress is yet to be made in understanding edit distribution and its implications.
@@ Appendix A Appendix
A.1 Implementation Details for ROME, MEMIT and EMMET
@@
A.2 Key-Value creation in ROME/MEMIT
@@
A.3 EMMET Derivation
[…]
As can be seen in the above equation, the preservation of representations happens in the first term whereas memorization of all the new facts are forced using an equality constraint in the second term. The above equation is solved using lagrange-multipliers. The Lagrangian for the above equation for multiple equality constraints requires a summation of lagrange multipliers and equals:
$\hat{W}K_{0}K_{0}^{T}=W_{0}K_{0}K_{0}^{T}+\sum_{i=1}^{E}\lambda_{i}k^{e^{T}}_{i}\$ (15)
which is same as:
$\hskip 8.0pt\hat{W}=W_{0}+\Lambda K_{E}^{T}C^{-1}_{0}\$ (17)
$\hat{W}=W_{0}+\Delta\hskip 10.0pt\text{where}$ (19)
A.4 EMMET - MEMIT Hyperparameter Comparison
Figures 9 - 16 present the comparison between EMMET and MEMIT for different hyperparameter values. The hyperparameter corresponds to the preservation term in the preservation memorization objective (equation 4). The figures show that both algorithm reach the same peak performance (Figure 9) across all models, but at different hyperparameter values. MEMIT reaches peak performance at lower hyperparameter values, whereas EMMET needs a larger weight for preservation to reach similar performance. This makes sense as EMMET works with a much stronger memorization constraint and thus requires larger weight to preserve the model by the same amount.
(a) GPT2-XL
(b) GPT-J
(c) Llama-2-7b
Figure 9: Comparison between EMMET and MEMIT for different hyperparameter values for the metric of Score. Hyperparameter controls the weight of preservation term over memorization term.
(a) GPT2-XL
(b) GPT-J
(c) Llama-2-7b
Figure 10: Comparison between EMMET and MEMIT for different hyperparameter values for the metric of Efficacy Score.
(a) GPT2-XL
(b) GPT-J
(c) Llama-2-7b
Figure 11: Comparison between EMMET and MEMIT for different hyperparameter values for the metric of Efficacy Magnitude.
(a) GPT2-XL
(b) GPT-J
(c) Llama-2-7b
Figure 12: Comparison between EMMET and MEMIT for different hyperparameter values for the metric of Paraphrase Score.
(a) GPT2-XL
(b) GPT-J
(c) Llama-2-7b
Figure 13: Comparison between EMMET and MEMIT for different hyperparameter values for the metric of Paraphrase Magnitude.
(a) GPT2-XL
(b) GPT-J
(c) Llama-2-7b
Figure 14: Comparison between EMMET and MEMIT for different hyperparameter values for the metric of Neighborhood Score.
(a) GPT2-XL
(b) GPT-J
(c) Llama-2-7b
Figure 15: Comparison between EMMET and MEMIT for different hyperparameter values for the metric of Neighborhood Magnitude.
(a) GPT2-XL
(b) GPT-J
(c) Llama-2-7b
Figure 16: Comparison between EMMET and MEMIT for different hyperparameter values for the metric of Generation Entropy.
A.5 EMMET and MEMIT Downstream Performance Comparison